# Offline Goal-Conditioned Reinforcement Learning with Projective Quasimetric Planning

### v2

@@ 4.3 Uniformly Mapping the Latent Space
with $\lambda_{\text{ood}}$ updated until satisfaction of a penalty $\big{(}(1-\delta)-\sum_{k=1}^{K}\psi(z_{k})\big{)}^{2}$, with $10<\delta<$ 1 the compliance. In other words, on average, all latent keypoints must stay in-distribution to some extent.
@@
Existence and Distinctness of Coulomb Energy Minimizers (See Appendix D.2.2) :Let’s consider a repulsive energy function $E:\mathcal{S}^{K}\rightarrow\mathbb{R}^{+}\cup\{+\infty\}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \{s_{k}\}_{k=1}^{K}\rightarrow\sum_{i\neq j}\frac{1}{d(s_{i},s_{j})}$. Under the assumptions of compactness of the state space $\mathcal{S}$ and continuity the quasimetric $d$ we have :
@@
(ii) For any such minimizer $\{s^{*}_{k}\}_{k=1}^{K}\subset\mathcal{S}\leavevmode\nobreak$, if $K>1$ then all keypoints are distinct.
@@ 4.4 Graph-based Navigation
Let’s consider $\mathcal{V}(g)=\{z_{1},\dots,z_{K},z_{g}\}$ be the latent key-points plus the current goal embedding. We keep an edge $(i,j)\in\mathcal{E}(g)$ only if $d(z_{i},z_{j})\leq\tau$ and set its weight to that distance. Here, $\tau$ is a stabilizing cut distance to avoid using potentially ill-approximated long distances. Consequently, we produce a directed graph $G(g)=(\mathcal{V}(g),\mathcal{E}(g))$, and all-pairs shortest paths are obtained with Floyd–Warshall algorithm floyd , yielding a matrix $D(g)^{*}$ that stores the distances from every keypoint to the goal within the learned graph.
@@
From a dataset transition $(s_{i},a_{i},s^{\prime}_{i})$ we sample a relatively short-range goal $g$ and compute the distance improvement $A_{i}$ in Equation 10.Advantage-Weighted Regression awr then fits a policy $\pi_{\theta_{\pi}}$ with the log-likelihood loss (11), weighted by $\exp(\alpha\cdot A_{i})$. We thus obtain a robust controller used to go from one keypoint to another.
@@
This regression objective has proven stable and effective for short-horizon offline control iql ; h_iql ; qphil .
@@ 5.1 Environments & Datasets
We evaluate on the OGBench suite ogbench , a recent benchmark tailored to Offline GCRL. We focus on the PointMaze environments, where a point-mass must reach goal locations across mazes of different scales (Figure 2) and a harder teleport variant with stochastic non-local transitions.
@@
We compare our proposed method to six baselines : Goal-Conditioned Behavioral Cloning (GCBC) gcbc , GC Implicit Value Learning (GCIVL) h_iql , GC Implicit Q Learning (GCIQL) iql , Contrastive RL (CRL) crl , Quasimetric RL (QRL) qrl , and Hierarchical Implicit Q Learning (HIQL) h_iql .
@@ 5.2.1 Performances
pointmaze-medium-navigate-v0 9$\pm$ 6 63$\pm$ 6 53$\pm$ 8 82$\pm$ 5 29$\pm$ 7 79$\pm$ 5 100$\pm$ 0
pointmaze-large-navigate-v0 29$\pm$ 6 45$\pm$ 5 34$\pm$ 3 86$\pm$ 9 39$\pm$ 7 58$\pm$ 5 99$\pm$ 1
pointmaze pointmaze-giant-navigate-v0 1$\pm$ 2 0$\pm$ 0 0$\pm$ 0 68$\pm$ 7 27$\pm$ 10 46$\pm$ 9 92$\pm$ 3
pointmaze-teleport-navigate-v0 25$\pm$ 3 45$\pm$ 3 24$\pm$ 7 4$\pm$ 4 24$\pm$ 6 18$\pm$ 4 43$\pm$ 0
pointmaze-medium-stitch-v0 23$\pm$ 18 70$\pm$ 14 21$\pm$ 9 80$\pm$ 12 0$\pm$ 1 74$\pm$ 6 99$\pm$ 1
pointmaze-large-stitch-v0 7$\pm$ 5 12$\pm$ 6 31$\pm$ 2 84$\pm$ 15 0$\pm$ 0 13$\pm$ 6 99$\pm$ 1
pointmaze-giant-stitch-v0 0$\pm$ 0 0$\pm$ 0 0$\pm$ 0 50$\pm$ 8 0$\pm$ 0 0$\pm$ 0 99$\pm$ 1
pointmaze-teleport-stitch-v0 31$\pm$ 9 44$\pm$ 2 25$\pm$ 3 9$\pm$ 5 4$\pm$ 3 34$\pm$ 4 35$\pm$ 7
Across all maze sizes and dataset types, ProQ consistently outperforms prior methods-particularly on the large and giant mazes. Its $92\%$ score on giant-navigate (versus $99\%$ score on giant-stitch) suggests that training on very long expert traces can introduce low-level control noise when learning. In the stochastic teleport variants ProQ’s is slightly below the best score. Moreover, ProQ’s lower standard deviations reflect a more stable latent mapping and planning pipeline.
@@ 5.2.3 Path Planning
Figure 4 depicts two sets of keypoints selected for different start and goal locations. The chosen keypoints trace the intuitive shortest corridors and never cut through walls, indicating that the learned quasimetric yields accurate one-step costs and that the offline Floyd–Warshall search exploits them effectively. Even on the longest route, only a dozen hops are required, demonstrating how a sparse, OOD-constrained graph can support near-optimal long-horizon navigation.
@@ B.1 Learning Agent & Mazes
OGBench is a standardized suite of offline goal-conditioned learning tasks. It includes PointMaze, a simple 2D navigation environment in which a point-mass agent (with continuous ($x,y$) position and bounded velocity inputs) must reach arbitrary goals within a bounded maze.
Four maze variants appear in OGBench (see Figure 6) : medium whose moderate trajectory lengths tests basic navigation ; large with longer corridor lengths increase the decision complexity ; giant scales the maze again, yielding an even larger layouts that force an agent to compose many short hops into a long-horizon path.teleport is of the same size as the large with stochastic portals that randomly chose exits (potentially a dead end).
@@ B.2 Datasets
Each layout provides two offline datasets of one million transitions. navigate contains long expert rollouts ($1000$ steps each) where a noisy expert wanders toward random goals ; these traces densely cover corridors but compound low-level control noise. stitch breaks those same trajectories into many shorter ones ($200$ steps each), forcing the agent to stitch discontiguous segments.
@@ C.1 Network architecture
If not otherwise stated, all our networks (state encoder, quasimetric embedder, odd detector, actor) are MLPs following the same structure : $3$ hidden layers of size $512$, GELU layer activation.Encoder $\phi$ maps a state $s\in\mathcal{S}$ into a latent space $\mathbb{R}^{16}$, and uses a MLP with residual connections. Quasimetric $d$ is an Interval Quasimetric Embedding (IQE) [66] with $64$ components, each of size $8$.OOD head $\psi$ is a scalar-output MLP with layer norm and sigmoid activation. Actor uses $0.1$dropout layers. It receives observations $s\in\mathcal{S}$ concatenated with target latent keypoints $z\in\mathbb{R}^{16}$ ; the outputs are a mean action $\mu(a|s,z)$ and a state–independent log-std for each action dimension.
@@ C.4 Data pipeline
At each gradient step, batches are sampled i.i.d. from the $1$ M-transition replay buffer considered. Goals for the value loss computation are drawn uniformly from the dataset ; while goals for the actor loss follow HER [2] trajectory-based geometric sampling, with a $\lambda=0.99$ parameter.
@@ D.1.1 Theoretical Analysis
In our learning framework, any generated state should be unreachable according to the classifier $\psi$, unless it matches a positive state (Section 4.2). Using these generated states, we learn $\psi$ with a binary cross-entropy loss with a dynamic Lagrange multiplier to enforce high confidence on positives while pushing negatives in-distribution probability towards 00 :
@@
  • Negative samples : $\tilde{z}=\phi(\tilde{s})$, $\tilde{s}\in\mathcal{\widetilde{D}}$, with the target $\psi(\tilde{z})=0$ ;
  • Loss : $\mathcal{L}_{\text{ood}}(z,\tilde{z})=-\lambda_{\text{ood}}\cdot\log\bigl{(}\psi(z)\bigr{)}-\log\leavevmode\nobreak\ \bigl{(}1-\psi(\tilde{z})\bigr{)}\ ,$
where $\lambda_{\text{ood}}$ is updated until satisfaction of a penalty $\Big{(}(1-\delta)-\psi(z)\Big{)}^{2}$, with $10<\delta<$ 1 the compliance.
@@
  • The dataset $\mathcal{D}\subset\mathcal{S}$, satisfies an $\epsilon$-density condition : $\exists\leavevmode\nobreak\ \epsilon>0,\forall\leavevmode\nobreak\ s\in\mathcal{S},\exists\leavevmode\nobreak\ s^{\prime}\in\mathcal{D}:\leavevmode\nobreak\ \lVert s-s^{\prime}\rVert_{2}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \epsilon\leavevmode\nobreak\ .$
  • We note as $Diam(\mathcal{D})$ the diameter of the dataset $\mathcal{D}$ : $Diam(\mathcal{D})=\max_{s,s^{\prime}\in\mathcal{D}}\leavevmode\nobreak\ \lVert s-s^{\prime}\rVert_{2}$
@@
    $\forall\leavevmode\nobreak\ s\in\mathbb{R}^{n}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \min_{s^{\prime}\in\mathcal{D}}\leavevmode\nobreak\ \lVert s-s^{\prime}\rVert_{2}\leq Diam(\mathcal{D})\Rightarrow\exists\leavevmode\nobreak\ \tilde{s}\in\mathcal{\widetilde{D}}:\leavevmode\nobreak\ \lVert s-\tilde{s}\rVert_{2}\leq\epsilon\leavevmode\nobreak\ .$ (12)
@@
    $\forall\leavevmode\nobreak\ s,s^{\prime}\in\mathbb{R}^{n}:\leavevmode\nobreak\ \lVert\phi(s)-\phi(s^{\prime})\rVert_{2}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ L_{\phi}\cdot\lVert s-s^{\prime}\rVert_{2}\leavevmode\nobreak\ .$ (13)
@@
    $\forall\leavevmode\nobreak\ z,z^{\prime}\in\mathbb{R}^{d}:\leavevmode\nobreak\ |\psi(z)-\phi(z^{\prime})|\leavevmode\nobreak\ \leq\leavevmode\nobreak\ L_{\psi}\cdot\lVert z-z^{\prime}\rVert_{2}\leavevmode\nobreak\ .$ (14)
  • After convergence of $\psi$, given the compliance $10<\delta<$ 1 , we have (Section 4.2) :
    $\forall\leavevmode\nobreak\ s\in\mathcal{D}:\psi(\phi(s))\geq 1-\delta\leavevmode\nobreak\ .$ (15)
@@
Concretely, in the assumptions above, the value $\bar{\delta}$ is the worst-case prediction of $\psi$ over all generated points (through interpolation or extrapolation) at least $\eta$ away from any state within the dataset $\mathcal{D}$.
@@
For any $s\in\mathcal{S}:\psi(\phi(s))\geq 1-\delta-(L_{\psi}L_{\phi})\epsilon$
@@
Let’s consider $s\in\mathcal{S}$. Since the dataset $\mathcal{D}$ is $\epsilon$-dense in $\mathcal{S}$, we know $\exists\leavevmode\nobreak\ s^{\prime}\in\mathcal{D},\leavevmode\nobreak\ \lVert s-s^{\prime}\rVert_{2}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \epsilon$, and by Lipschitz continuity :
@@
In particular : $\psi(\phi(s))-\psi(\phi(s^{\prime}))\geq-(L_{\psi}L_{\phi})\epsilon\leavevmode\nobreak$. By considering this inequality and Equation 15 :
$\psi(\phi(s))\geq\psi(\phi(s^{\prime}))-(L_{\psi}L_{\phi})\epsilon\geq 1-\delta-(L_{\psi}L_{\phi})\epsilon\leavevmode\nobreak\ .$
@@
For any $s\in\mathbb{R}^{n}$ : $\eta\leq\min_{s^{\prime}\in\mathcal{D}}\lVert s-s^{\prime}\rVert_{2}\leq Diam(\mathcal{D})\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \psi(\phi(s))\leq\bar{\delta}+(L_{\psi}L_{\phi})\epsilon\leavevmode\nobreak\ .$
@@
Let’s consider $s\in\mathbb{R}^{n}$ with $\rho=min_{s^{\prime}\in\mathcal{D}}\lVert s-s^{\prime}\rVert$ that verifies $\eta\leq\rho\leq Diam(\mathcal{D})$. Since the set of generated states $\mathcal{\widetilde{D}}$ is $\epsilon$-dense in a shell around $\mathcal{D}$, up to a distance $Diam(\mathcal{D}$, we know $\exists\leavevmode\nobreak\ s^{\prime}\in\mathcal{\widetilde{D}},\leavevmode\nobreak\ \lVert s-s^{\prime}\rVert_{2}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \epsilon$, and by Lipschitz continuity :
@@
In particular : $\psi(\phi(s))-\psi(\phi(s^{\prime}))\leq(L_{\psi}L_{\phi})\epsilon\leavevmode\nobreak$. By considering this inequality and Equation 16 :
$\psi(\phi(s))\leq\psi(\phi(s^{\prime}))+(L_{\psi}L_{\phi})\epsilon\leq\bar{\delta}+(L_{\psi}L_{\phi})\epsilon\leavevmode\nobreak\ .$

### v3

@@ 4.3 Uniformly Mapping the Latent Space
with $\lambda_{\text{ood}}$ updated until satisfaction of a penalty $\big((1-\delta)-\sum_{k=1}^{K}\psi(z_{k})\big)^{2}$, with $0<\delta<1$ the compliance. In other words, on average, all latent keypoints must stay in-distribution to some extent.
@@
Existence and Distinctness of Coulomb Energy Minimizers (See Appendix D.2.2) :Let’s consider a repulsive energy function $E:\mathcal{S}^{K}\rightarrow\mathbb{R}^{+}\cup\{+\infty\}\penalty 10000\ ,\penalty 10000\ \{s_{k}\}_{k=1}^{K}\rightarrow\sum_{i\neq j}\frac{1}{d(s_{i},s_{j})}$. Under the assumptions of compactness of the state space $\mathcal{S}$ and continuity the quasimetric $d$ we have :
@@
(ii) For any such minimizer $\{s^{*}_{k}\}_{k=1}^{K}\subset\mathcal{S}\penalty 10000\$, if $K>1$ then all keypoints are distinct.
@@ 4.4 Graph-based Navigation
Let’s consider $\mathcal{V}(g)=\{z_{1},\dots,z_{K},z_{g}\}$ be the latent key-points plus the current goal embedding. We keep an edge $(i,j)\in\mathcal{E}(g)$ only if $d(z_{i},z_{j})\leq\tau$ and set its weight to that distance. Here, $\tau$ is a stabilizing cut distance to avoid using potentially ill-approximated long distances. Consequently, we produce a directed graph $G(g)=(\mathcal{V}(g),\mathcal{E}(g))$, and all-pairs shortest paths are obtained with Floyd–Warshall algorithm Weisstein (2008), yielding a matrix $D(g)^{*}$ that stores the distances from every keypoint to the goal within the learned graph.
@@
From a dataset transition $(s_{i},a_{i},s^{\prime}_{i})$ we sample a relatively short-range goal $g$ and compute the distance improvement $A_{i}$ in Equation 10.Advantage-Weighted Regression Peng et al. (2019) then fits a policy $\pi_{\theta_{\pi}}$ with the log-likelihood loss (11), weighted by $\exp(\alpha\cdot A_{i})$. We thus obtain a robust controller used to go from one keypoint to another.
@@
This regression objective has proven stable and effective for short-horizon offline control Kostrikov et al. (2021); Park et al. (2023); Canesse et al. (2024).
@@ 5.1 Environments & Datasets
We evaluate on the OGBench suite Park et al. (2024a), a recent benchmark tailored to Offline GCRL. We focus on the PointMaze environments, where a point-mass must reach goal locations across mazes of different scales (Figure 2) and a harder teleport variant with stochastic non-local transitions.
@@
We compare our proposed method to six baselines : Goal-Conditioned Behavioral Cloning (GCBC) Ding et al. (2019), GC Implicit Value Learning (GCIVL) Park et al. (2023), GC Implicit Q Learning (GCIQL) Kostrikov et al. (2021), Contrastive RL (CRL) Eysenbach et al. (2023), Quasimetric RL (QRL) Wang et al. (2023), and Hierarchical Implicit Q Learning (HIQL) Park et al. (2023).
@@ 5.2.1 Performances
[…]
Across all maze sizes and dataset types, ProQ consistently outperforms prior methods-particularly on the large and giant mazes. Its $92\%$ score on giant-navigate (versus $99\%$ score on giant-stitch) suggests that training on very long expert traces can introduce low-level control noise when learning. In the stochastic teleport variants ProQ’s is slightly below the best score. Moreover, ProQ’s lower standard deviations reflect a more stable learning pipeline.
@@ 5.2.3 Path Planning
Figure 4 depicts two sets of keypoints selected for different start and goal locations. The chosen keypoints trace the intuitive shortest corridors and never cut through walls, indicating that the learned quasimetric yields accurate one-step costs and that the offline Floyd–Warshall search exploits them effectively. Even on the longest route, only a few hops are required, demonstrating how a sparse graph can support near-optimal long-horizon navigation.
@@ 6 Discussion
7 Acknowledgement
This project was provided with computer and storage resources by GENCI at IDRIS thanks to the grant 2024-AD011015210 on the supercomputer Jean Zay’s CSL partition.
@@ B.1 Learning Agent & Mazes
OGBench is a standardized suite of offline goal‐conditioned learning tasks. It includes PointMaze, a simple 2D navigation environment in which a point‐mass agent (with continuous ($x,y$) position and bounded velocity inputs) must reach arbitrary goals within a bounded maze.
Four maze variants appear in OGBench (see Figure 6) : medium whose moderate trajectory lengths tests basic navigation ; large with longer corridor lengths increase the decision complexity ; giant scales the maze again, yielding an even larger layouts that force an agent to compose many short hops into a long‐horizon path.teleport is of the same size as the large with stochastic portals that randomly chose exits (potentially a dead end).
@@ B.2 Datasets
Each layout provides two offline datasets of one million transitions. navigate contains long expert rollouts ($1000$ steps each) where a noisy expert wanders toward random goals ; these traces densely cover corridors but compound low‐level control noise. stitch breaks those same trajectories into many shorter ones ($200$ steps each), forcing the agent to stitch discontiguous segments.
@@ C.1 Network architecture
If not otherwise stated, all our networks (state encoder, quasimetric embedder, odd detector, actor) are MLPs following the same structure : $3$ hidden layers of size $512$, GELU layer activation.Encoder $\phi$ maps a state $s\in\mathcal{S}$ into a latent space $\mathbb{R}^{16}$, and uses a MLP with residual connections. Quasimetric $d$ is an Interval Quasimetric Embedding (IQE) [66] with $64$ components, each of size $8$.OOD head $\psi$ is a scalar‐output MLP with layer norm and sigmoid activation. Actor uses 0.10.1dropout layers. It receives observations $s\in\mathcal{S}$ concatenated with target latent keypoints $z\in\mathbb{R}^{16}$ ; the outputs are a mean action $\mu(a|s,z)$ and a state–independent log‐std for each action dimension.
@@ C.4 Data pipeline
At each gradient step, batches are sampled i.i.d. from the $1$ M‐transition replay buffer considered. Goals for the value loss computation are drawn uniformly from the dataset ; while goals for the actor loss follow HER [2] trajectory‐based geometric sampling, with a $\lambda=0.99$ parameter.
@@ D.1.1 Theoretical Analysis
In our learning framework, any generated state should be unreachable according to the classifier $\psi$, unless it matches a positive state (Section 4.2). Using these generated states, we learn $\psi$ with a binary cross-entropy loss with a dynamic Lagrange multiplier to enforce high confidence on positives while pushing negatives in-distribution probability towards 0 :
@@
[…]
where $\lambda_{\text{ood}}$ is updated until satisfaction of a penalty $\Big((1-\delta)-\psi(z)\Big)^{2}$, with $0<\delta<1$ the compliance.
@@
  • The dataset $\mathcal{D}\subset\mathcal{S}$, satisfies an $\epsilon$-density condition : $\exists\penalty 10000\ \epsilon>0,\forall\penalty 10000\ s\in\mathcal{S},\exists\penalty 10000\ s^{\prime}\in\mathcal{D}:\penalty 10000\ \lVert s-s^{\prime}\rVert_{2}\penalty 10000\ \leq\penalty 10000\ \epsilon\penalty 10000\ .$
  • We note as $Diam(\mathcal{D})$ the diameter of the dataset $\mathcal{D}$ : $Diam(\mathcal{D})=\max_{s,s^{\prime}\in\mathcal{D}}\penalty 10000\ \lVert s-s^{\prime}\rVert_{2}$ .
@@
    $\forall\penalty 10000\ s\in\mathbb{R}^{n}\penalty 10000\ ,\penalty 10000\ \min_{s^{\prime}\in\mathcal{D}}\penalty 10000\ \lVert s-s^{\prime}\rVert_{2}\leq Diam(\mathcal{D})\Rightarrow\exists\penalty 10000\ \tilde{s}\in\mathcal{\widetilde{D}}:\penalty 10000\ \lVert s-\tilde{s}\rVert_{2}\leq\epsilon\penalty 10000\ .$ (12)
@@
    $\forall\penalty 10000\ s,s^{\prime}\in\mathbb{R}^{n}:\penalty 10000\ \lVert\phi(s)-\phi(s^{\prime})\rVert_{2}\penalty 10000\ \leq\penalty 10000\ L_{\phi}\cdot\lVert s-s^{\prime}\rVert_{2}\penalty 10000\ .$ (13)
@@
    $\forall\penalty 10000\ z,z^{\prime}\in\mathbb{R}^{d}:\penalty 10000\ |\psi(z)-\phi(z^{\prime})|\penalty 10000\ \leq\penalty 10000\ L_{\psi}\cdot\lVert z-z^{\prime}\rVert_{2}\penalty 10000\ .$ (14)
  • After convergence of $\psi$, given the compliance $0<\delta<1$ , we have (Section 4.2) :
    $\forall\penalty 10000\ s\in\mathcal{D}:\psi(\phi(s))\geq 1-\delta\penalty 10000\ .$ (15)
@@
Concretely, in the assumptions above, the value $\bar{\delta}$ is the worst‐case prediction of $\psi$ over all generated points (through interpolation or extrapolation) at least $\eta$ away from any state within the dataset $\mathcal{D}$.
@@
For any $s\in\mathcal{S}:\psi(\phi(s))\geq 1-\delta-(L_{\psi}L_{\phi})\epsilon$ .
@@
Let’s consider $s\in\mathcal{S}$. Since the dataset $\mathcal{D}$ is $\epsilon$-dense in $\mathcal{S}$, we know $\exists\penalty 10000\ s^{\prime}\in\mathcal{D},\penalty 10000\ \lVert s-s^{\prime}\rVert_{2}\penalty 10000\ \leq\penalty 10000\ \epsilon$, and by Lipschitz continuity :
@@
In particular : $\psi(\phi(s))-\psi(\phi(s^{\prime}))\geq-(L_{\psi}L_{\phi})\epsilon\penalty 10000\$. By considering this inequality and Equation 15 :
$\psi(\phi(s))\geq\psi(\phi(s^{\prime}))-(L_{\psi}L_{\phi})\epsilon\geq 1-\delta-(L_{\psi}L_{\phi})\epsilon\penalty 10000\ .$
@@
For any $s\in\mathbb{R}^{n}$ : $\eta\leq\min_{s^{\prime}\in\mathcal{D}}\lVert s-s^{\prime}\rVert_{2}\leq Diam(\mathcal{D})\penalty 10000\ \Rightarrow\penalty 10000\ \psi(\phi(s))\leq\bar{\delta}+(L_{\psi}L_{\phi})\epsilon\penalty 10000\ .$
@@
Let’s consider $s\in\mathbb{R}^{n}$ with $\rho=min_{s^{\prime}\in\mathcal{D}}\lVert s-s^{\prime}\rVert$ that verifies $\eta\leq\rho\leq Diam(\mathcal{D})$. Since the set of generated states $\mathcal{\widetilde{D}}$ is $\epsilon$-dense in a shell around $\mathcal{D}$, up to a distance $Diam(\mathcal{D}$, we know $\exists\penalty 10000\ s^{\prime}\in\mathcal{\widetilde{D}},\penalty 10000\ \lVert s-s^{\prime}\rVert_{2}\penalty 10000\ \leq\penalty 10000\ \epsilon$, and by Lipschitz continuity :
@@
In particular : $\psi(\phi(s))-\psi(\phi(s^{\prime}))\leq(L_{\psi}L_{\phi})\epsilon\penalty 10000\$. By considering this inequality and Equation 16 :
$\psi(\phi(s))\leq\psi(\phi(s^{\prime}))+(L_{\psi}L_{\phi})\epsilon\leq\bar{\delta}+(L_{\psi}L_{\phi})\epsilon\penalty 10000\ .$